Betancourt International has operations in Arrakis. The balance sheet for this division in Arrakeen solaris shows assets of 40,000 solaris, debt in the amount of 12,500 solaris, and equity of 27,500 solaris.

a.

If the current exchange ratio is 1.35 solaris per dollar, what does the balance sheet look like in dollars? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)


b.

Assume that one year from now the balance sheet in solaris is exactly the same as at the beginning of the year. If the exchange rate is 1.45 solaris per dollar, what does the balance sheet look like in dollars now? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)


c.

Assume that one year from now the balance sheet in solaris is exactly the same as at the beginning of the year. If the exchange rate is 1.26 solaris per dollar, what does the balance sheet look like in dollars now? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

a). Assets = solaris 40,000 / ($ / solaris 1.35) = $29,629.63

Debt = solaris 12,500 / ($ / solaris 1.35) = $9,259.27

Equity = solaris 27,500 / ($ / solaris 1.35) = $20,370.37

Assets Amount Liabilities Amount
Debt $9,259.27
Equity $20.370.37
Assets $29,629.64 Total Liabilities & Equity $29,629.64

b). Assets = solaris 40,000 / ($ / solaris 1.45) = $27,586.21

Debt = solaris 12,500 / ($ / solaris 1.45) = $8,620.69

Equity = solaris 27,500 / ($ / solaris 1.45) = $18,965.52

Assets Amount Liabilities Amount
Debt $8,620.69
Equity $18,965.52
Assets $27,586.21 Total Liabilities & Equity $27,586.21

c). Assets = solaris 40,000 / ($ / solaris 1.26) = $31,746.03

Debt = solaris 12,500 / ($ / solaris 1.26) = $9,920.63

Equity = solaris 27,500 / ($ / solaris 1.26) = $21,825.40

Assets Amount Liabilities Amount
Debt $9,920.63
Equity $21,825.40
Assets $31,746.03 Total Liabilities & Equity $31,746.03